OpenStudy (sissyedgar):

On Monday, a clothing store sold 25 pairs of pants. On Tuesday, the store sold 37 pairs of pants. What was the percent of increase in number of pants sold from Monday to Tuesday?

OpenStudy (sissyedgar):

@Holly00d1248

OpenStudy (sissyedgar):

@paki

OpenStudy (paki):

you know the percentage increase or decrease formula...?

OpenStudy (sissyedgar):

i dont get any of this!

OpenStudy (paki):

rite.. .find the difference between the percentages... multiply that difference with 100... then divide it on the ist percentage.. .

OpenStudy (sissyedgar):

i got 48... but how to i but tht 2 a %?

OpenStudy (paki):

37-25 = 12.... 12*100 = 1200... divide by 25 = 1200/25 = 48.. and that is the answe.r...

OpenStudy (sissyedgar):

i need it in a percentage!

OpenStudy (sissyedgar):

-_-

OpenStudy (sissyedgar):

where u go @paki?

OpenStudy (paki):

48 is percentage...

OpenStudy (sissyedgar):

SO 48%... duh me! :P srry

OpenStudy (paki):

yeah 48 %...